OpenStudy (anonymous):

Find d2y/dx2 in terms of x and y. y6 = x9

OpenStudy (anonymous):

y^6 and X^9

OpenStudy (anonymous):

are you supposed to find the second derivative of \[y^6=x^9\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok here goes. the derivative of \[x^9\] is \[9x^8\] and the derivative of \[y^6\] is \[6y^5y'\] so first step is to write \[6y^5y'=9x^8\]

OpenStudy (anonymous):

solve for \[y'\] to get \[y'=\frac{3x^8}{2y^5}\]

OpenStudy (anonymous):

clear so far, because that was the easy part

OpenStudy (anonymous):

now you have to find \[y''\] using the quotient rule

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so the quotient rule: the denominator will be \[4y^{10}\]

OpenStudy (anonymous):

the numerator will be \[2y^5\times 24x^7-3x^8\times 10y^4y'\] is that apart clear?